CDMS 2012: A DATABASE WITHIN THE FRAMEWORK OF THE VIRTUAL ATOMIC AND MOLECULAR DATA CENTRE − RECENT DEVELOPMENTS

\underline{HOLGER S. P. MÜLLER}, C. P. ENDRES, J. STUTZKI, S. SCHLEMMER, I. Physikalisches Institut, Universität zu Köln, 50937 Köln, Germany.

The catalog section of the Cologne Database for Molecular Spectroscopy, CDMS,^{a}^{b}^{c} contains experimental and calculated transition frequencies of molecules and atoms observable by radioastronomical means with information on uncertainties, intensities and quantum numbers. Initially constructed as ascii tables, its inclusion into a database environment within the framework of the European Framework Programme 7 project Virtual Atomic and Molecular Data Centre (VAMDC)^{d} has begun in June 2008. The conversion required time because of the extensive content of the CDMS, the desire to maintain the existing functionality, the additional requisites as part of the VAMDC framework, and the sustained development of the VAMDC infrastructure. The new CDMS is scheduled to be available publicly prior to this conference; the old CDMS will be accessible for a transitional period. We will present some of the new CDMS features, provide some overview on new or updated entries from our last account three years ago^{e} as well as activities in other sections of the CDMS.

aInternet address: http://www.astro.uni-koeln.de/cdms/

bH. S. P. Müller, F. Schlöder, J. Stutzki, and G. Winnewisser, J. Mol. Struct. 742, 215−227 (2005)

cH. S. P. Müller, S. Thorwirth, D. A. Roth, and G. Winnewisser, Astron. Astrophys. 370, L49−L52 (2001)

dInternet address: http://www.vamdc.org/

eH. S. P. Müller, J. Stutzki, S. Schlemmer, contribution WH07, presented at the 64th International Symposium on Molecular Spectroscopy, Columbus, OH, USA, 2009